4. Valuing semiannual coupon bonds Aa Aa Bonds often pay a coupon twice a year. For the valuation of bonds that make semiannual payments, the number of periods doubles, whereas the amount of cash flow decreases by half. Using the values of cash flows and number of periods, the valuation model is adjusted accordingly Assume that a $1,000,000 par value, semiannual coupon US Treasury note with three years to maturity (YTM) has a coupon rate of 6%. The yield to maturity of the bond is 11.00%. Using this information and ignoring the other costs involved, calculate the value of the Treasury note O $551,320.40 O $743,844.98 O $1,050,134.09 O $875,111.74 Based on your calculations and understanding of semiannual coupon bonds, complete the following statement: Assuming that interest rates remain constant, the T-notes price is expected toPlease help and explain
Answer #1

Answer 1:

Correct option is:

$875,111.74

Explanation:

Par value = $1,000,000

Semiannual coupon amount = $1,000,000 * 6% / 2 = $30,000

Years to maturity = 3 years

Number of semiannual periods = 3 * 2 = 6

Semiannual yield to maturity = 11%/2 = 5.5%

To get value of Treasury note, we will use PV function of Excel:

PV ( rate, nper, pmt, fv, type)

PV (5.5%, 6, 30000, 1000000, 0)

= $875,111.74

As such option D is correct and other options A, B, and C are incorrect.

Answer 2:

Assuming that interest rates remain constant, the T - note's price is expected to increase.

Explanation:

Assuming that interest rates remain constant, as the time to maturity decreases, T - note's price is expected to increase.